OpenStudy (anonymous):

(^3√5x) (^4√2x) I got ^4√10x and it only gave me partial credit so i did something wrong..?

OpenStudy (phi):

is this \[\sqrt[3]{5x}\cdot \sqrt[4]{2x}\] because this does not really simplify. On the other hand if it is \[\sqrt[4]{5x}\cdot \sqrt[4]{2x}\] you can rewrite it as \[\sqrt[4]{10x^2}\]

OpenStudy (anonymous):